The sum of a number and twice its square is 105. Find the number.
Answer by mananth(16949)

let the number be x
x+2x^2=105
2x^2+x-105 =0
2x^2+15x-14x-105
x(2x+15)-7(2x+15)=0
(x-7)(2x+15)=0
x= 7 OR x=-7.5